OpenStudy (anonymous):

solve for b: (remember: b^2 equals b to the second power-"exponent") b^2 = 8b + 84

OpenStudy (welshfella):

b^2 -8b - 84 = 0 b^2 - 14b + 6b - 84 = 0 now factor by grouping

OpenStudy (anonymous):

b(b-14)+6(b-14)............= (b-14)(b+6) (b-14)=0..........................or((b+6)=0 b=14...............or b= -6